# Exactly solvable Schrödinger operators related to the hypergeometric equation

### v1

@@ 2.1 Hypergeometric equation
The hypergeometric equation is closely related to the so-called Riemann equation, that is the class of equations on the Riemann sphere $\mathbb{C}$$\cup\{\infty\}$ having 3 regular singular points. Suppose these points are $z_{1},z_{2},z_{3}$, and the corresponding indices are $\rho_{1},{\tilde{\rho}}_{1}$, resp. $\rho_{2}$, ${\tilde{\rho}}_{2}$, resp. $\rho_{3}$, ${\tilde{\rho}}_{3}$. Then we have the constraint
$\rho_{1}+{\tilde{\rho}}_{1}+\rho_{2}+{\tilde{\rho}}_{2}+\rho_{3}+{\tilde{\rho}}_{3}=1.$ (2.3)
Without limiting the generality we can put $z_{3}$ at $\infty$. The corresponding Riemann equation is given by the Riemann operator
${\cal P}\left(\begin{array}[]{cccc}z_{1}&z_{2}&\infty&\\ \rho_{1}&\rho_{2}&\rho_{3}&z,\partial_{z}\\ {\tilde{\rho}}_{1}&{\tilde{\rho}}_{2}&{\tilde{\rho}}_{3}&\end{array}\right)$ (2.8)
$:=$ $\partial_{z}^{2}-\left(\frac{\rho_{1}+{\tilde{\rho}}_{1}-1}{z-z_{1}}+\frac{\rho_{2}+{\tilde{\rho}}_{2}-1}{z-z_{2}}\right)\partial_{z}$
$+\frac{\rho_{1}{\tilde{\rho}}_{1}(z_{1}-z_{2})}{(z-z_{1})^{2}(z-z_{2})}+\frac{\rho_{2}{\tilde{\rho}}_{2}(z_{2}-z_{1})}{(z-z_{2})^{2}(z-z_{1})}+\frac{\rho_{3}{\tilde{\rho}}_{3}}{(z-z_{1})(z-z_{2})}.$
Here is the relation between the hypergeometric operator and the Riemann operator:
$\begin{array}[]{rl}{\cal F}(a,b;c;z,\partial_{z})&=z(1-z){\cal P}\left(\begin{array}[]{cccc}0&1&\infty&\\ 0&0&a&z,\partial_{z}\\ 1-c&c-a-b&b&\end{array}\right)\\[19.91692pt] &=z(1-z)\partial_{z}^{2}+(c-(a+b+1)z)\partial_{z}-ab.\end{array}$ (2.14)
@@
$\begin{array}[]{rrr}\alpha=c-1&\ \ \beta=a+b-c,&\ \ \mu=a-b;\\ a=\frac{1+\alpha+\beta-\mu}{2},&\ \ b=\frac{1+\alpha+\beta+\mu}{2},&\ \ c=1+\alpha;\end{array}$ (2.17)
@@ 2.2 Gegenbauer equation
Here is its relationship to the Riemann operator:
@@
$\mathbf{Z}_{\alpha,\lambda}(w)$ $=\frac{(w\pm 1)^{-\frac{1}{2}-\alpha-\lambda}}{\Gamma(\lambda+1)}F\Big{(}\frac{1}{2}+\alpha+\lambda,\frac{1}{2}+\lambda;2\lambda+1;\frac{2}{1\pm w}\Big{)}.$ (2.22)
@@ 2.3 Liouville transformation
$\left(p(z)\partial_{z}^{2}+q(z)\partial_{z}+r(z)\right)u(z),$ (2.24)
@@ 3 Gegenbauer Hamiltonians
We have reinterpreted the Gegenbauer equation as the eigenequation of a certain operator $L_{\alpha}$ with the eigenvalue $\lambda^{2}$. It is natural to interpret this operator as acting on functions on an interval contained in $\mathbb{C}$}, whose endpoints are singularities of the Gegenbauer equation. In each of these cases we perform the Liouville transformation, which yields a 1-dimensional Hamiltonian. We will consider three cases:
@@
Figure 1: Gegenbauer equation on the $w$ plane. The hyperbolic Hamiltonian acts on the interval marked with a dotted line, the spherical Hamiltonian—with a thick line and the deSitterian Hamiltonian—with a dashed line.
@@ 3.1 Spherical case
Proof.  $\mathcal{P}^{\rm s}_{\alpha,\lambda}(r)$ and $\mathcal{P}^{\rm s}_{-\alpha,\lambda}(r)$ can be used as a basis of solutions of (3.9). The connection formula found using (A.23) is
@@
For ${\rm Re}~\alpha>-1$ the integral kernel (3.16) is square integrable, and hence it defines a Hilbert-Schmidt operator. It depends analytically on $\alpha$. Hence it defines a holomorphic family of operators. For ${\rm Re}~\alpha\geq 1$, $r^{\frac{1}{2}-\alpha}$ is not $L^{2}$-integrable. Hence for such $\alpha$ (3.10) possesses a unique closed realization.
The singularities of the Gamma function yield the discrete spectrum. $\Box$
@@ 3.2 Hyperbolic case
$\sigma_{\rm d}(L_{\alpha}^{\rm h})$ $=\emptyset,\quad-\frac{1}{2}<{\rm Re}\,\alpha;$ (3.32)
@@
For ${\rm Re}\,\alpha>-1$ and ${\rm Re}\,\lambda>0$ the functions $\mathcal{P}^{\rm h}_{\alpha,\lambda}(r)$ resp. $\mathcal{Q}^{\rm h}_{\alpha,\lambda}(r)$, are square integrable at the endpoints. We check that under these conditions the integral kernel 3.34 defines a bounded operator, depending analytically on $\alpha$.
@@ 3.3 DeSitterian case
$\mathcal{Q}^{\mathrm{dS}}_{\alpha,\lambda}(r):={\rm e}^{-{\rm i}\frac{\pi}{2}(\frac{1}{2}+\alpha+\lambda)}\frac{(\cosh r)^{\alpha+\frac{1}{2}}}{2^{\lambda}}~{\cal Z}_{\alpha,\lambda}(-{\rm i}\sinh r).$ (3.41)
@@
We check that the integral kernel 3.44 defines a bounded operator.
@@
The Whipple transformation (A.26), (A.27) on the interval $w\in]-1,1[$ has two versions: above this interval and below, that is
@@ 4 Hypergeometric Hamiltonians of the first kind
The hypergeometric equation has been reinterpreted as the eigenequation of the operator $L_{\alpha,\beta}$ with the eigenvalue $\frac{\mu^{2}}{4}$. It is natural to interpret this operator as acting on functions on an interval whose endpoints are singularities of the hypergeometric equations. In each of these cases we perform the Liouville transformation, which yields a 1-dimensional Hamiltonian. We will consider three cases:
@@
Figure 2: Hypergeometirc Hamiltonians of the first kind on the $z$-plane. The hyperbolic Hamiltonian act on the interval with the dotted line, spherical Hamiltonian—the thick line, and deSitterian Hamiltonian—the dashed line.
@@ 4.1 Spherical case
$P_{\alpha,\beta,\mu}^{\rm s}(r)$ $:=\Big{(}\sin\frac{r}{2}\Big{)}^{\alpha+\frac{1}{2}}\Big{(}\cos\frac{r}{2}\Big{)}^{\beta+\frac{1}{2}}\mathbf{F}\left(\frac{\alpha+\beta+\mu+1}{2},\frac{\alpha+\beta-\mu+1}{2};1+\alpha;\sin^{2}\Big{(}\frac{r}{2}\Big{)}\right)$ (4.14)
@@
Proof.  Considering $P^{\rm s}_{\alpha,\beta,\mu}(r)$, and $P^{\rm s}_{-\alpha,\beta,\mu}(r)$ as a basis of solutions of (4.11), we can rewrite the connection formula (A.9) as
@@ 4.2 Hyperbolic case
$P_{\alpha,\beta,\mu}^{\rm h}(r)$ $:=\Big{(}\sinh\frac{r}{2}\Big{)}^{\alpha+\frac{1}{2}}\Big{(}\cosh\frac{r}{2}\Big{)}^{\beta+\frac{1}{2}}\mathbf{F}\left(\frac{\alpha+\beta+\mu+1}{2},\frac{\alpha+\beta-\mu+1}{2};1+\alpha;-\sinh^{2}\frac{r}{2}\right)$ (4.28)
@@
$Q_{\alpha,\beta,\mu}^{\rm h}(r)$ $:=\Big{(}\sinh\frac{r}{2}\Big{)}^{-\mu-\beta-\frac{1}{2}}\Big{(}\cosh\frac{r}{2}\Big{)}^{\beta+\frac{1}{2}}\mathbf{F}\left(\frac{\alpha+\beta+\mu+1}{2},\frac{-\alpha+\beta+\mu+1}{2};1+\mu;-\sinh^{-2}\frac{r}{2}\right)$ (4.30)
@@
$\sigma_{\rm d}(L_{\alpha,\beta}^{\rm h})=$ $\left\{-\Big{(}\frac{\alpha+\beta}{2}+k\Big{)}^{2}\quad|\quad k\in{\mathbb{N}}_{0}+\frac{1}{2},\quad k<-{\rm Re}\,\frac{\alpha+\beta}{2}\right\},$ (4.37)
@@
Proof.  Considering $P^{\rm h}_{-\alpha,\beta,\mu}(r)$, and $P^{\rm h}_{-\alpha,\beta,\mu}(r)$ as a basis of solutions of (4.49), we can rewrite connection formula (A.13) as
@@
${\cal W}\left(Q^{\rm h}_{\alpha,\beta,\mu}(r),P^{\rm h}_{\alpha,\beta,\mu}(r)\right)$ $=\frac{1}{\Gamma\left(\frac{1+\alpha+\beta+\mu}{2}\right)\Gamma\left(\frac{1+\alpha-\beta+\mu}{2}\right)}.$
@@ 4.3 DeSitterian case
$\sigma_{\rm d}(L_{\alpha.\beta}^{\mathrm{dS}})=$ $\left\{-\left(\frac{\alpha+\beta}{2}-k\right)^{2}\bigg{|}\quad k\in{\mathbb{N}}_{0}+\frac{1}{2},\quad k<{\rm Re}\,\frac{\alpha+\beta}{2}\right\}$ (4.59)
$\cup$ $\left\{-\left(\frac{\alpha-\beta}{2}-k\right)^{2}\bigg{|}\quad k\in{\mathbb{N}}_{0}+\frac{1}{2},\quad k<{\rm Re}\frac{\alpha-\beta}{2}\right\},$
@@
Proof.  In the connection formula (A.15) we insert $z=\frac{1-{\rm i}s}{2}$ and multiply it by $\Big{(}\frac{1-{\rm i}s}{2}\Big{)}^{\frac{\alpha}{2}+\frac{1}{4}}\Big{(}\frac{1+{\rm i}s}{2}\Big{)}^{\frac{\beta}{2}+\frac{1}{4}}$, obtaining
@@
Using (A.12), we know that
@@ 5 Hypergeometric Hamiltonians of the second kind
We have reinterpreted (5.5) as the eigenequation of the operator $K_{\kappa,\mu}$ with eigenvalue $\delta$.
Again, it is natural to interpret this operator as acting on functions on an interval whose endpoints are singularities of the hypergeometric equations. In each of these cases we perform the Liouville transformation, which yields a 1-dimensional Hamiltonian. We will consider three cases:
@@
Figure 3: Hypergeometric Hamiltonian of the second kind on the $z$-plane. The hyperbolic Hamiltonian act on the interval with dotted line, the spherical—with the thick line, and the Desitterian—with the dashed line.
@@ 5.1 Spherical case
$\mathbb{Q}_{\alpha,\beta,\mu}^{\rm s}(u)$ (5.14)
$:=$ $\left(\frac{{\rm i}}{1-{\rm e}^{2{\rm i}u}}\right)^{\frac{-1-\beta-\mu}{2}}\left(\frac{-{\rm i}}{1-{\rm e}^{-2{\rm i}u}}\right)^{\frac{\beta}{2}}\mathbf{F}\Big{(}\frac{\alpha+\beta+\mu+1}{2},\frac{-\alpha+\beta+\mu+1}{2};\mu+1;1-{\rm e}^{2{\rm i}u}\Big{)}$ (5.15)
$=$ $\left(\frac{{\rm i}}{1-{\rm e}^{2{\rm i}u}}\right)^{\frac{\alpha}{2}}\left(\frac{-{\rm i}}{1-{\rm e}^{-2{\rm i}u}}\right)^{\frac{-1-\alpha-\mu}{2}}\mathbf{F}\Big{(}\frac{\alpha+\beta+\mu+1}{2},\frac{\alpha-\beta+\mu+1}{2};\mu+1;1-{\rm e}^{-2{\rm i}u}\Big{)}.$ (5.16)
@@
$\frac{1}{(K^{\rm s}_{\tau,\mu}-\delta)}(x,y)=$ $\frac{\Gamma\left(\frac{1-\alpha-\beta+\mu}{2}\right)\Gamma\left(\frac{1+\alpha-\beta+\mu}{2}\right)\Gamma\left(\frac{1-\alpha+\beta+\mu}{2}\right)\Gamma\left(\frac{1+\alpha+\beta+\mu}{2}\right)}{4\pi}$
@@
${\cal W}\left(\mathbb{Q}^{\rm s}_{\beta,\alpha,\mu}(\pi-u),\mathbb{Q}^{\rm s}_{\alpha,\beta,\mu}(u)\right)$ $=2{\cal W}\left(Q^{\mathrm{dS}}_{\beta,\alpha,\mu}(-r),Q^{\mathrm{dS}}_{\alpha,\beta,\mu}(r)\right)$ (5.28)
@@
${\cal W}\big{(}\mathbb{Q}^{\rm s}_{\beta,\alpha,-\mu}(\pi-u),\mathbb{Q}^{\rm s}_{\alpha,\beta,\mu}(u)\big{)}$ $=2{\cal W}\left(Q^{\mathrm{dS}}_{\beta,\alpha,-\mu}(-r),Q^{\mathrm{dS}}_{\alpha,\beta,\mu}(r)\right)$ (5.30)
$=\frac{2({\rm e}^{{\rm i}\pi\frac{\mu}{2}}\cos\pi\beta+{\rm e}^{-{\rm i}\pi\frac{\mu}{2}}\cos\pi\alpha)}{\pi}.$ (5.31)
The $L^{2}$ integrability condition for $\mathbb{Q}^{\rm s}_{\alpha,\beta,\mu}(u)$ at 0 and $\mathbb{Q}^{s}_{\beta,\alpha,\mu}(\pi-u)$ at $\pi$ is ${\rm Re}\,\mu>-2$. The $L^{2}$ norm of this kernel 5.25 is finite. For ${\rm Re}(\mu)\geq 2$ it is a unique candidate for the resolvent.
@@
For $\mu\in{\mathbb{Z}}$ we have an additional identity for the $\mathbb{Q}^{\rm s}$ function, which follows directly from (A.17):

### v2

@@ 2.1 Hypergeometric equation
$\begin{array}[]{rrr}\alpha=c-1&\ \ \beta=a+b-c,&\ \ \mu=a-b;\\ a=\frac{1+\alpha+\beta+\mu}{2},&\ \ b=\frac{1+\alpha+\beta-\mu}{2},&\ \ c=1+\alpha;\end{array}$ (2.5)
@@
To understand the properties and importance of the hypergeometric equation it is useful to view it as a special case of the Riemann equation–the most general equation on the Riemann sphere possessing at most 3 singular points, all of them regular-singular. There are many books and survey articles that comprehensively treat the Riemann and hypergeometric equations, including [WW, Ince, NU, NIST1]. Throughout this paper, we follow the notation established in [DW, D1, D2]. For the convenience of the reader in Appendix B we explain how the hypergeometric equation arises from the Riemann equation.
@@ 2.2 Gegenbauer equation
Here is its relationship to the Riemann operator (defined in (B.68)):
@@
$\mathbf{Z}_{\alpha,\lambda}(w)$ $=\frac{(w\pm 1)^{-\frac{1}{2}-\alpha-\lambda}}{\Gamma(\lambda+1)}F\left(\frac{1}{2}+\alpha+\lambda,\frac{1}{2}+\lambda;2\lambda+1;\frac{2}{1\pm w}\right).$ (2.10)
@@ 2.3 Liouville transformation
$\left(p(z)\partial_{z}^{2}+q(z)\partial_{z}+r(z)\right)u(z){=0},$ (2.12)
@@ 3 Gegenbauer Hamiltonians
We have interpreted the Gegenbauer equation as the eigenequation of the operator $L_{\alpha}$ with the eigenvalue $\lambda^{2}$ acting on functions on the complex plane. We would like to reinterpret it as an operator on functions of the real variable $r\in]a,b[$ for some $\infty\leq a<b\leq+\infty$. To this end we need to insert the intervals ]a,b[]a,b[ into the complex plane. We will consider three intervals; in each case their endpoints are among the singular points of the Gegenbauer equation (that is, $1,-1,\infty$). For each of these intervals we perform the Liouville transformation, which yields a 1-dimensional Hamiltonian on $C_{\mathrm{c}}^{\infty}]a,b[$. As we will see below, in each of these cases we will obtain a natural holomorphic family of closed realizations possessing elegant formulas for the Green functions.
Here are the three cases that we consider:
@@
Figure 1: Gegenbauer equation on the $w$ plane.
The spherical Hamiltonian acts on the interval marked with a thick line, the hyperbolic Hamiltonian—with a dotted line, and the deSitterian Hamiltonian—with a dashed line.
@@ 3.1 Spherical case
Proof. Treating $\mathcal{P}^{\rm s}_{\alpha,\lambda}(r)$ and $\mathcal{P}^{\rm s}_{-\alpha,\lambda}(r)$ as a basis of solutions of (3.9), we obtain a connection formula found using (C.22):
@@
Thus (3.16) is an integral kernel of the form (D.7). Let us denote it by $R_{\alpha}^{\rm s}(-\lambda^{2};x,y)$.
Let us fix temporarily $\lambda$. For ${\rm Re}\,\alpha>-1$ the integral kernel $R_{\alpha}^{\rm s}(-\lambda^{2};x,y)$ is square integrable, and hence it defines a Hilbert-Schmidt operator, which we denote $R_{\alpha}^{\rm s}(-\lambda^{2})$. By Appendix D.2 there exists an operator $L_{\alpha}^{\rm s}$, which is a closed realization of the left hand side of (3.10) such that $R_{\alpha}^{\rm s}(-\lambda^{2})=(L_{\alpha}^{\rm s}+\lambda^{2})^{-1}$.
The Hilbert-Schmidt norm of $R_{\alpha}^{\rm s}(-\lambda^{2})$ is uniformly bounded for $\alpha\in K$ with $K\subset\{{\rm Re}\,\alpha>-1\}$ compact. Besides, $R_{\alpha}^{\rm s}(-\lambda^{2};x,y)$ for any $x,y$ depends analytically on $\alpha$. Hence for any $f,g\in C_{\mathrm{c}}(]0,\pi[)$
$\alpha\mapsto\int\overline{f(x)}R_{\alpha}^{\rm s}(-\lambda^{2};x,y)g(y){\rm d}x{\rm d}y$ (3.22)
is analytic. Using Thm A.1 we obtain that $\{{\rm Re}\,\alpha>0\}\ni\alpha\mapsto R_{\alpha}^{\rm s}(-\lambda^{2})\in B\big(L^{2}]0,\pi[\big)$ is an analytic family of bounded operators. Hence, by Thm A.2, $\{{\rm Re}\,\alpha>0\}\ni\alpha\mapsto L_{\alpha}^{\rm s}\in C\big(L^{2}]0,\pi[\big)$ is an analytic family of closed operators.
For ${\rm Re}\,\alpha\geq 1$, $r^{\frac{1}{2}-\alpha}$ is not $L^{2}$-integrable, and only $r^{\frac{1}{2}+\alpha}$ is. Hence for such $\alpha$ the right hand side of (3.10) possesses a unique closed realization, which has to coincide with the operator $L_{\alpha}^{\rm s}$ that we have just constructed. In particular, $L_{\alpha}^{\rm s}$ does not depend on $\lambda$. By the uniqueness of analytic continuation, $L_{\alpha}^{\rm s}$ does not depend on the choice of $\lambda$ for all ${\rm Re}\,\alpha>-1$.
The singularities of the Gamma function yield the discrete spectrum of $L_{\alpha}^{\rm s}$. $\Box$
@@ 3.2 Hyperbolic case
$\sigma_{\rm d}(L_{\alpha}^{\rm h})$ $=\emptyset,\quad-\frac{1}{2}{\leq}{\rm Re}\,\alpha;$ (3.33)
@@
For ${\rm Re}\,\alpha>-1$ and ${\rm Re}\,\lambda>0$ the functions $\mathcal{P}^{\rm h}_{\alpha,\lambda}(r)$ resp. $\mathcal{Q}^{\rm h}_{\alpha,\lambda}(r)$, are square integrable at the endpoints. Using the Schur test we check that under these conditions the integral kernel 3.35 defines a bounded operator, depending analytically on $\alpha$.
@@ 3.3 DeSitterian case
$\mathcal{Q}^{\mathrm{dS}}_{\alpha,\lambda}(r):={\rm e}^{-{\rm i}\frac{\pi}{2}(\frac{1}{2}+\alpha+\lambda)}\frac{(\cosh r)^{\alpha+\frac{1}{2}}}{2^{\lambda}}~{\mathbf{Z}_{\alpha,\lambda}}(-{\rm i}\sinh r).$ (3.42)
@@
Using the Schur test we check that the integral kernel (3.45) defines a bounded operator.
@@
The Whipple transformation (C.25), (C.26) on the interval $w\in]-1,1[$ has two versions: above this interval and below, that is
@@ 4 Hypergeometric Hamiltonians of the first kind
The hypergeometric equation has been interpreted as the eigenequation of the operator $L_{\alpha,\beta}$ with the eigenvalue $\frac{\mu^{2}}{4}$ acting on functions on the complex plane. Let us reinterpret this operator as acting on functions on three intervals whose endpoints are singularities of the hypergeometric equations. We will see again that these choices lead to simple formulas for Green functions. In each of these cases we perform the Liouville transformation, which yields a 1-dimensional Hamiltonian. We will consider three cases:
@@
Figure 2: Hypergeometric Hamiltonians of the first kind on the $z$-plane.
The spherical Hamiltonian acts on the interval marked with a thick line, the hyperbolic Hamiltonian—with a dotted line, and deSitterian Hamiltonian—with a dashed line.
@@ 4.1 Spherical case
$P_{\alpha,\beta,\mu}^{\rm s}(r)$ $:=\Big(\sin\frac{r}{2}\Big)^{\alpha+\frac{1}{2}}\Big(\cos\frac{r}{2}\Big)^{\beta+\frac{1}{2}}\mathbf{F}\left(\frac{\alpha+\beta+\mu+1}{2},\frac{\alpha+\beta-\mu+1}{2};1+\alpha;\sin^{2}\Big(\frac{r}{2}\Big)\right)$
@@
Proof. Considering $P^{\rm s}_{\alpha,\beta,\mu}(r)$, and $P^{\rm s}_{-\alpha,\beta,\mu}(r)$ as a basis of solutions of (4.11), we can rewrite the connection formula (C.9) as
@@ 4.2 Hyperbolic case
$P_{\alpha,\beta,\mu}^{\rm h}(r)$ $:=\Big(\sinh\frac{r}{2}\Big)^{\alpha+\frac{1}{2}}\Big(\cosh\frac{r}{2}\Big)^{\beta+\frac{1}{2}}\mathbf{F}\left(\frac{\alpha+\beta+\mu+1}{2},\frac{\alpha+\beta-\mu+1}{2};1+\alpha;-\sinh^{2}\frac{r}{2}\right)$
@@
$Q_{\alpha,\beta,\mu}^{\rm h}(r)$ $:=\Big(\sinh\frac{r}{2}\Big)^{-\mu-\beta-\frac{1}{2}}\Big(\cosh\frac{r}{2}\Big)^{\beta+\frac{1}{2}}\mathbf{F}\left(\frac{\alpha+\beta+\mu+1}{2},\frac{-\alpha+\beta+\mu+1}{2};1+\mu;-\sinh^{-2}\frac{r}{2}\right)$
@@
$\sigma_{\rm d}(L_{\alpha,\beta}^{\rm h})=$ $\left\{-\Big(\frac{\alpha+\beta}{2}+k\Big)^{2}\quad|\quad k\in{\mathbb{N}}_{0}+\frac{1}{2},\quad k<-{\rm Re}\,\frac{\alpha+\beta}{2}\right\},$
@@
Proof. Considering $P^{\rm h}_{-\alpha,\beta,\mu}(r)$, and $P^{\rm h}_{-\alpha,\beta,\mu}(r)$ as a basis of solutions of (4.46), we can rewrite connection formula (C.13) as
@@
${\cal W}\left(Q^{\rm h}_{\alpha,\beta,\mu}(r),P^{\rm h}_{\alpha,\beta,\mu}(r)\right)$ $=\frac{1}{\Gamma\left(\frac{1+\alpha+\beta+\mu}{2}\right)\Gamma\left(\frac{1+\alpha-\beta+\mu}{2}\right)}.$ (4.39)
@@ 4.3 DeSitterian case
$\sigma_{\rm d}(L_{\alpha.\beta}^{\mathrm{dS}})=$ $\left\{-\left(\frac{\alpha+\beta}{2}-k\right)^{2}\bigg|\quad k\in{\mathbb{N}}_{0}+\frac{1}{2},\quad k<{\rm Re}\,\frac{\alpha+\beta}{2}\right\}$
$\cup$ $\left\{-\left(\frac{\alpha-\beta}{2}-k\right)^{2}\bigg|\quad k\in{\mathbb{N}}_{0}+\frac{1}{2},\quad k<{\rm Re}\frac{\alpha-\beta}{2}\right\},$ (4.56)
@@
Proof. In the connection formula (C.15) we insert $z=\frac{1-{\rm i}s}{2}$ and multiply it by $\Big(\frac{1-{\rm i}s}{2}\Big)^{\frac{\alpha}{2}+\frac{1}{4}}\Big(\frac{1+{\rm i}s}{2}\Big)^{\frac{\beta}{2}+\frac{1}{4}}$, obtaining
@@
Using (C.12), we know that
@@ 5 Hypergeometric Hamiltonians of the second kind
We have interpreted (5.5) as the eigenequation of the operator $K_{\kappa,\mu}$ with eigenvalue $\delta$. Again, we will reinterpret this operator as acting on functions on an interval whose endpoints are singularities of the hypergeometric equations. In each of these cases we perform the Liouville transformation, which yields a 1-dimensional Hamiltonian. We will consider three cases:
@@
Figure 3: Hypergeometric Hamiltonian of the second kind on the $z$-plane.
The spherical Hamiltonian acts on the interval marked with a thick line, the hyperbolic—with a dotted line, and the DeSitterian—with a dashed line.
@@ 5.1 Spherical case
$\mathbb{Q}_{\alpha,\beta,\mu}^{\rm s}(u)$
$:=$ $\left(\frac{{\rm i}}{1-{\rm e}^{2{\rm i}u}}\right)^{\frac{-1-\beta-\mu}{2}}\left(\frac{-{\rm i}}{1-{\rm e}^{-2{\rm i}u}}\right)^{\frac{\beta}{2}}\mathbf{F}\left(\frac{\alpha+\beta+\mu+1}{2},\frac{-\alpha+\beta+\mu+1}{2};\mu+1;1-{\rm e}^{2{\rm i}u}\right)$
$=$ $\left(\frac{{\rm i}}{1-{\rm e}^{2{\rm i}u}}\right)^{\frac{\alpha}{2}}\left(\frac{-{\rm i}}{1-{\rm e}^{-2{\rm i}u}}\right)^{\frac{-1-\alpha-\mu}{2}}\mathbf{F}\left(\frac{\alpha+\beta+\mu+1}{2},\frac{\alpha-\beta+\mu+1}{2};\mu+1;1-{\rm e}^{-2{\rm i}u}\right).$ (5.14)
@@
$\frac{1}{\left(K^{\rm s}_{\tau,\mu}-\delta\right)}(x,y)=$ $\frac{\Gamma\left(\frac{1-\alpha-\beta+\mu}{2}\right)\Gamma\left(\frac{1+\alpha-\beta+\mu}{2}\right)\Gamma\left(\frac{1-\alpha+\beta+\mu}{2}\right)\Gamma\left(\frac{1+\alpha+\beta+\mu}{2}\right)}{4\pi}$
@@
${\cal W}\left(\mathbb{Q}^{\rm s}_{\beta,\alpha,\mu}(\pi-u),\mathbb{Q}^{\rm s}_{\alpha,\beta,\mu}(u)\right)$ $=2{\cal W}\left(Q^{\mathrm{dS}}_{\beta,\alpha,\mu}(-r),Q^{\mathrm{dS}}_{\alpha,\beta,\mu}(r)\right)$
@@
${\cal W}\big(\mathbb{Q}^{\rm s}_{\beta,\alpha,-\mu}(\pi-u),\mathbb{Q}^{\rm s}_{\alpha,\beta,\mu}(u)\big)$ $=2{\cal W}\left(Q^{\mathrm{dS}}_{\beta,\alpha,-\mu}(-r),Q^{\mathrm{dS}}_{\alpha,\beta,\mu}(r)\right)$
$=\frac{2\left({\rm e}^{{\rm i}\pi\frac{\mu}{2}}\cos\pi\beta+{\rm e}^{-{\rm i}\pi\frac{\mu}{2}}\cos\pi\alpha\right)}{\pi}.$ (5.27)
The $L^{2}$ integrability condition for $\mathbb{Q}^{\rm s}_{\alpha,\beta,\mu}(u)$ at 0 and $\mathbb{Q}^{s}_{\beta,\alpha,\mu}(\pi-u)$ at $\pi$ is ${\rm Re}\,\mu>-2$. The $L^{2}$ norm of this kernel 5.23 is finite. For ${\rm Re}(\mu)\geq 2$ it is a unique candidate for the resolvent.
@@
For $\mu\in{\mathbb{Z}}$ we have an additional identity for the $\mathbb{Q}^{\rm s}$ function, which follows directly from (C.17):